The students of same age group from two different schools were compared for variability in their mathematical skill. A random sample of 25 pupils from one school had a variance of 16 marks while a random sample of 22 pupils from the other school had a variance of 8 marks. Examine if the difference in variability is significant. [F at (24,21) at 5% level of significance is 2.05]



1
Expert's answer
2021-12-10T15:15:19-0500

Solution

We are given:

"Sample\:size\:\left(n_1\right)=25"

"Sample\:size\:\left(n_2\right)=22"

"S_1^2=16"

"S_2^2=8"

Now we have to compute:

"H_o:\sigma_1^2=\sigma_2^2"

"H_o:\sigma_1^2 \ne" "\sigma_2^2"

The test statistic is as follows:

"F=\frac{16}{8}" "=2"

"F_{\frac{0.5}{2},\:24,\:21}=2.05"

Here "F<F_{\frac{0.5}{2},\:24,\:21}"

Thus, "H_o" is accepted.We can say that their is no variability in mathematical skills in one school of other school.